### v2

# A robust parameterized enhanced shift-splitting preconditioner for three-by-three block saddle point problems

Abstract
This paper proposes a new parameterized enhanced shift-splitting (PESS) preconditioner to solve the three-by-three block saddle point problem (SPP). Additionally, we introduce a local PESS (LPESS) preconditioner by relaxing the PESS preconditioner. Necessary and sufficient criteria are established for the convergence of the proposed PESS iterative process for any random initial guess. Furthermore, we meticulously investigate the spectral bounds of the PESS and LPESS preconditioned matrices. Moreover, empirical investigations have been performed for the sensitivity analysis of the proposed PESS preconditioner, which unveils its robustness. Numerical experiments are carried out to demonstrate the enhanced efficiency and robustness of the proposed PESS and LPESS preconditioners compared to the existing block diagonal and shift-splitting preconditioners.
Keywords. Krylov subspace iterative process $\cdot$ Preconditioner $\cdot$ Shift-splitting $\cdot$ Saddle point problem $\cdot$ Robust $\cdot$ GMRES
AMS subject classification. 65F08 $\cdot$ 65F10 $\cdot$ 65F50
1 Introduction
Three-by-three block saddle point problems (SPPs) have recently received considerable attention due to their wide applicability in various computational research and technological applications. These domains include the constraint optimization problems [16], computational fluid dynamics [8, 14], the constrained least square estimations [6], optimal control [20], solving Maxwell equations [12], magma-mantle dynamics [21] and so on. The three-by-three block SPP is represented as follows [18]:
$\mathcal{A}{\bf u}={\bf d},$ (1.1)
where $\mathcal{A}=\left[\begin{matrix}A&B^{T}&0\\ -B&0&-C^{T}\\ 0&C&0\end{matrix}\right],\,{\bf u}=\left[\begin{matrix}x\\ y\\ z\end{matrix}\right],\,{\bf d}=\left[\begin{matrix}f\\ g\\ h\end{matrix}\right],$ and $A\in{\mathbb{R}}^{n\times n}$ is a symmetric positive definite (SPD) matrix, $B\in{\mathbb{R}}^{m\times n}$ and $C\in{\mathbb{R}}^{p\times m}$ are the full row rank matrices. Here, $B^{T}$ stands for the transpose of $B,$ and $f\in{\mathbb{R}}^{n},\,g\in{\mathbb{R}}^{m}$ and $h\in{\mathbb{R}}^{p}$ are known vectors. Under the stated assumptions on block matrices A, B, and C, $\mathcal{A}$ is nonsingular, which ensures the existence of a unique solution for the system (1.1).
Due to the growing popularity of three-by-three block SPPs in a variety of applications, it is essential to determine the efficient and robust numerical solution of three-by-three block SPPs (1.1). However, the indefiniteness and poor spectral properties of $\mathcal{A}$ makes challenging to determine the solution of the system (1.1). Furthermore, $\mathcal{A}$ is generally sparse and large; therefore, iterative processes become superior to direct methods. The Krylov subspace iterative process is preferable among other iterative processes investigated widely in the literature due to their minimal storage requirement and feasible implementation [30]. Slow convergence of the Krylov subspace iterative process is a major drawback for system of linear equations of large dimensions. Moreover, the saddle point matrix $\mathcal{A}$ can be very sensitive, which leads to a significant slowdown in the solution algorithm. Therefore, it is important to develop novel, efficient, and robust preconditioners for the fast convergence of the Krylov subspace iterative process, which can handle the sensitivity of the problem (1.1) when the matrix $\mathcal{A}$ is ill-conditioned.
Consider the following partitioning of $\mathcal{A}:$
[…]
then the linear system (1.1) can be identified as the standard two-by-two block SPP [5]. In the last decades, colossal attention has been given to the development of numerical solution methods for standard two-by-two block SPP, such as Uzawa methods [7, 13, 2], Krylov subspace iterative process [17, 26], null space methods [15, 25]. However, the properties of the submatrices in (1.8) are different from the standard two-by-two block SPP. Notice that the leading blocks are skew-symmetric and positive semidefinite in the first partitioning. The second partitioning highlights that the $(1,2)$ block is rank deficient, and the $(2,2)$ block is skew-symmetric. In contrast, in the standard SPP, $(1,2)$ block is full row rank, and the $(2,2)$ block is generally symmetric positive semidefinite. Therefore, the existing literature for solving the two-by-two SPPs (see [5]) may not be applied to solve (1.1) directly, and it is essential to develop new preconditioners for three-by-three block SPP that exploit the specific structure of the saddle point matrix $\mathcal{A}$, which is sensitive in nature.
Recently, various effective preconditioners have been developed in the literature for solving three-by-three block SPPs. For instance, Huang and Ma [18] has studied the exact block diagonal (BD) and inexact BD (IBD) preconditioners ${\mathscr{P}}_{BD}$ and ${\mathscr{P}}_{IBD}$ in the following forms:
${\mathscr{P}}_{BD}=\left[\begin{matrix}A&0&0\\ 0&S&0\\ 0&0&CS^{-1}C^{T}\end{matrix}\right]\quad\mbox{and}\quad{\mathscr{P}}_{IBD}=\left[\begin{matrix}\widehat{A}&0&0\\ 0&\widehat{S}&0\\ 0&0&C\widehat{S}^{-1}C^{T}\end{matrix}\right],$ (1.9)
where $S=BA^{-1}B^{T},$ $\widehat{A}$ and $\widehat{S}$ are SPD approximations of $A$ and $S,$ respectively. Although the spectrum of the preconditioned matrices ${\mathscr{P}}^{-1}_{BD}\mathcal{A}$ and ${\mathscr{P}}^{-1}_{IBD}\mathcal{A}$ have good clustering properties, these preconditioners have certain shortcomings, such as they are time-consuming, expensive, require excessive number of iterations, and condition numbers are very large. For more details, see the paper [18]. To overcome these drawbacks of BD preconditioners, researchers have started exploring shift-splitting (SS) preconditioners, which was initially developed for a non-Hermitian system of linear equations by Bai et al. [3] and later for two-by-two block SPPs [10, 23, 11]. Recently, Cao [9] enhanced this idea and introduced the following SS preconditioner ${\mathscr{P}}_{SS}$ and relaxed SS (RSS) preconditioner ${\mathscr{P}}_{RSS}$ for the three-by-three block SPP (1.1)
${\mathscr{P}}_{SS}=\frac{1}{2}\left[\begin{matrix}\alpha I+A&B^{T}&0\\ -B&\alpha I&-C^{T}\\ 0&C&\alpha I\end{matrix}\right]\,\,\mbox{and}\,\,{\mathscr{P}}_{RSS}=\frac{1}{2}\left[\begin{matrix}A&B^{T}&0\\ -B&\alpha I&-C^{T}\\ 0&C&\alpha I\end{matrix}\right],$ (1.10)
where $\alpha>0$ and verified unconditionally convergence of the associated SS iterative process. Here, $I$ stands for the identity matrix of the appropriate dimension. Wang and Zhang [27] generalized the SS preconditioner by introducing a new parameter $\beta>0$ in the $(3,3)$ block. Further, Yin et al. [31] developed an extensive generalized SS (EGSS) preconditioner
${\mathscr{P}}_{EGSS}=\frac{1}{2}\left[\begin{matrix}\alpha P+A&B^{T}&0\\ -B&\beta Q&-C^{T}\\ 0&C&\gamma W\end{matrix}\right],$ (1.11)
for the three-by-three block SPP (1.1), where $\alpha,\beta,\gamma>0$ and $P,Q,W$ are SPD matrices and investigated its convergent properties. There are a number of other preconditioners and iterative processes to solve the three-by-three block SPP (1.1) has been studied recently in the literature [19, 1, 24, 29, 4].
Despite exhibiting favorable performance of SS, RSS and EGSS preconditioners, they do not outperform BD and IBD preconditioners. Therefore, to improve the convergence speed and efficiency of the preconditioners ${\mathscr{P}}_{SS},{\mathscr{P}}_{RSS}$ and ${\mathscr{P}}_{EGSS},$ this paper presents a parameterized enhanced shift-splitting (PESS) preconditioner by introducing a parameter $s>0$ for there-by-three block SPPs. As per the direct correlation between the rate of convergence in Krylov subspace iterative processes and the spectral properties of the preconditioned matrix, we perform in-depth spectral distribution of PESS preconditioned matrix. We summarize the following main contributions of this paper:
  • We propose the PESS iterative process along with its associated PESS preconditioner and its relaxed version known as the local PESS (LPESS) preconditioner to solve the three-by-three block SPP (1.1). The proposed preconditioners generalize the existing SS preconditioners, for example, [9, 27, 31].
  • General framework is given on necessary and sufficient criteria for the convergence of the PESS iterative process. These investigations also encompass the unconditional convergence of other exiting SS preconditioners in the literature [9, 31].
  • We have conducted a comprehensive analysis of the spectral distribution for the PESS preconditioned matrix. Our framework allows us to derive the spectral distribution for the SS and EGSS preconditioned matrices.
  • We empirically show that our proposed preconditioner significantly reduces the condition number of the ill-conditioned saddle point matrix $\mathcal{A}$. Thereby establishing an efficiently solvable, well-conditioned system.
  • Numerical experiments show that the proposed PESS and LPESS preconditioners outperform the existing BD and SS preconditioners. Moreover, we perform sensitivity analysis by perturbing the ill-conditioned matrix $\mathcal{A}$ in (1.1), demonstrating the robustness of the proposed preconditioner.
The structure of this article is as follows. In Section 2, we propose the PESS iterative process and the associated PESS preconditioner. Section 3 is devoted to investigating the convergence of the PESS iterative process. In Section 4, we investigate the spectral distribution of the PESS preconditioned matrix. In Section 5, we present the LPESS preconditioner. In Section 6, we conduct numerical experiments to illustrate the computational efficiency and robustness of the proposed preconditioner. In the end, a brief concluding remark is provided in Section 7.
Notation. The following notations are adopted throughout the paper. For any matrix $A\in{\mathbb{R}}^{n\times n},$ $A^{T},\sigma(A)$ and $\vartheta(A)$ denotes transpose, spectrum and spectral radius of $A,$ respectively. For any $x\in\mathbb{C}^{n},$ $x^{*}$ denotes its conjugate transpose. The symbol $\mathrm{diag}(A,B)$ represents a block diagonal matrix with diagonal blocks $A$ and $B.$ The real and imaginary parts of any complex number $z$ are represented by $\Re(z)$ and $\Im(z)$, respectively. For any symmetric matrix $A\in{\mathbb{R}}^{n\times n},$ $\lambda_{\max}(A)$ and $\lambda_{\min}(A)$ denotes the maximum and minimum eigenvalue of $A,$ respectively. The notation $A>0$ signifies that $A$ is SPD, and $A>B$ identifies $A-B>0.$ Abbreviations used in the paper are as follows:
SPP: saddle point problem, SPD: symmetric positive definite, BD: block diagonal, IBD: inexact block diagonal, SS: shift-splitting, RSS: relaxed shift-splitting, EGSS: extensive generalized shift-spliting, PESS: parameterized enhanced shift-splitting, LPESS: local parameterized enhanced shift-splitting GMRES: generalized minimum residual process.
2 The parameterized enhanced shift-splitting (PESS) iterative process and the preconditioner
In this section, we proposed a parameterized enhanced shift-splitting (PESS) iterative process and the corresponding PESS preconditioner. Let $s>0$ be a real number. Then, we split the matrix $\mathcal{A}$ in the form
$\mathcal{A}=(\Sigma+s\mathcal{A})-(\Sigma-(1-s)\mathcal{A})=:{\mathscr{P}}_{PESS}-\mathcal{Q}_{PESS},\,\mbox{where}$ (2.1)
${\mathscr{P}}_{PESS}=\left[\begin{matrix}\Lambda_{1}+sA&sB^{T}&0\\ -sB&\Lambda_{2}&-sC^{T}\\ 0&sC&\Lambda_{3}\end{matrix}\right],\mathcal{Q}_{PESS}=\left[\begin{matrix}\Lambda_{1}-(1-s)A&-(1-s)B^{T}&0\\ (1-s)B&\Lambda_{2}&(1-s)C^{T}\\ 0&-(1-s)C&\Lambda_{3}\end{matrix}\right],$
$\Sigma=\mathrm{diag}(\Lambda_{1},\Lambda_{2},\Lambda_{3}),$ and $\Lambda_{1},\Lambda_{2},\Lambda_{3}$ are SPD matrices. The matrix ${\mathscr{P}}_{PESS}$ is nonsingular for $s>0.$ The special matrix splitting (2.1) leads us to the subsequent iterative process for solving the three-by-three block SPP given in (1.1):
The PESS iterative process. Let $s$ be a positive constant, and let $\Lambda_{1}\in{\mathbb{R}}^{n\times n},\Lambda_{2}\in{\mathbb{R}}^{m\times m}$ $\Lambda_{3}\in{\mathbb{R}}^{p\times p}$ be SPD matrices. For any initial guess ${\bf u}_{0}\in{\mathbb{R}}^{n+m+p}$ and $k=0,1,2\ldots,$ until a specified termination criterion is fulfilled, we compute
${\bf u}_{k+1}=\mathcal{T}{\bf u}_{k}+{\bf c},$ (2.2)
where ${\bf u}_{k}=[x_{k}^{T},y_{k}^{T},z_{k}^{T}]^{T},{\bf c}={\mathscr{P}}_{PESS}^{-1}{\bf d},$ and $\mathcal{T}={\mathscr{P}}_{PESS}^{-1}\mathcal{Q}_{PESS}$ is called the iteration matrix for the PESS iterative process.
Moreover, the matrix splitting (2.2) introduces a new preconditioner ${\mathscr{P}}_{PESS}$, identified as PESS preconditioner. This preconditioner is a generalization of the previous work in the literature, see for example [9, 27, 31] for specific choices of $\Lambda_{1},\Lambda_{2},\Lambda_{3}$ and $s,$ which are listed in Table 1.
Table 1: ${\mathscr{P}}_{PESS}$ as a generalization of the above SS preconditioners for different choices of $\Lambda_{1},\Lambda_{2},\Lambda_{3}$ and $s.$
${\mathscr{P}}_{PESS}$ $\Lambda_{1}$ $\Lambda_{2}$ $\Lambda_{3}$ $s$ memo
${\mathscr{P}}_{SS}$ [9] $\frac{1}{2}\alpha I$ $\frac{1}{2}\alpha I$ $\frac{1}{2}\alpha I$ $s=\frac{1}{2}$ $\alpha>0$
${\mathscr{P}}_{GSS}$ [27] $\frac{1}{2}\alpha I$ $\frac{1}{2}\alpha I$ $\frac{1}{2}\beta I$ $s=\frac{1}{2}$ $\alpha,\beta>0$
${\mathscr{P}}_{EGSS}$ [31] $\frac{1}{2}\alpha P$ $\frac{1}{2}\beta Q$ $\frac{1}{2}\gamma W$ $s=\frac{1}{2}$ $P,Q,W$ are SPD
matrices and $\alpha,\beta,\gamma>0$
In the implementation of the PESS iterative process or the PESS preconditioner to enhance the rate of convergence of the Krylov subspace iterative process like GMRES, at each iterative step, we solve the following system of linear equations
${\mathscr{P}}_{PESS}w=r,$ (2.3)
where $r=[r_{1}^{T},r_{2}^{T},r_{3}^{T}]^{T}\in{\mathbb{R}}^{n+m+p}$ and $w=[w_{1}^{T},w_{2}^{T},w_{3}^{T}]^{T}\in{\mathbb{R}}^{n+m+p}.$ Specifying $\widehat{X}=\Lambda_{2}+s^{2}C^{T}\Lambda_{3}^{-1}C$ and $\widehat{A}=\Lambda_{1}+sA+s^{2}B^{T}\widehat{X}^{-1}B,$ then ${\mathscr{P}}_{PESS}$ can be written in the following way
${\mathscr{P}}_{PESS}=\left[\begin{matrix}I&sB^{T}\widehat{X}^{-1}&0\\ 0&I&-sC^{T}\Lambda_{3}^{-1}\\ 0&0&I\end{matrix}\right]\left[\begin{matrix}\widehat{A}&0&0\\ 0&\widehat{X}&0\\ 0&0&\Lambda_{3}\end{matrix}\right]\left[\begin{matrix}I&0&0\\ -s\widehat{X}^{-1}B&I&0\\ 0&s\Lambda_{3}^{-1}C&I\end{matrix}\right].$ (2.4)
The decomposition (2.4) leads us to the following Algorithm 1 to determine the solution of the system (2.3).
Algorithm 1 Computation of $w$ from ${\mathscr{P}}_{PESS}w=r$
1:Solve $\widehat{X}v_{1}=r_{2}+sC^{T}\Lambda_{3}^{-1}r_{3}$ to find $v_{1};$
2:Compute $v=r_{1}-sB^{T}v_{1};$
3:Solve $\widehat{A}w_{1}=v$ to find $w_{1}$
4:Solve $\widehat{X}v_{2}=sBw_{1}$ to find $v_{2}$;
5:Compute $w_{2}=v_{1}+v_{2};$
6:Solve $\Lambda_{3}w_{3}=r_{3}-sCw_{2}$ for $w_{3}.$
The implementation of Algorithm 1 requires to solve two linear subsystems with coefficient matrices $\widehat{X}$ and $\widehat{A}$. Since $\Lambda_{1},\Lambda_{2},\Lambda_{3},\widehat{X},$ and $\widehat{A}$ are SPD and $s>0$, we can apply the Cholesky factorization to solve these linear subsystems exactly or inexactly by the preconditioned conjugate gradient method.
3 Convergence analysis of the PESS iterative process
In this section, we investigate the convergence of the PESS iterative process (2.2). To achieve this aim, the following definition and lemmas are crucial.
Definition 3.1.
A matrix $\mathcal{A}$ is called positive stable if $\Re(\lambda)>0$ for all $\lambda\in\sigma(\mathcal{A})$.
Lemma 3.1.
[9] Let $A\in{\mathbb{R}}^{n\times n}$ be a SPD matrix, and let $B\in{\mathbb{R}}^{m\times n}$ and $C\in{\mathbb{R}}^{p\times m}$ be full row rank matrices. Then, the saddle point matrix $\mathcal{A}$ is positive stable.
Lemma 3.2.
Let $A\in{\mathbb{R}}^{n\times n}$ be a SPD matrix, and let $B\in{\mathbb{R}}^{m\times n}$ and $C\in{\mathbb{R}}^{p\times m}$ be full row rank matrices. Then $\Sigma^{-1}\mathcal{A}$ ((((or, $\Sigma^{-\frac{1}{2}}\mathcal{A}\Sigma^{-\frac{1}{2}})$ is positive stable.
Proof.
Since $\Sigma^{-\frac{1}{2}}=\mathrm{diag}(\Lambda_{1}^{-\frac{1}{2}},\Lambda_{2}^{-\frac{1}{2}},\Lambda_{3}^{-\frac{1}{2}}),$ we obtain
$\Sigma^{-\frac{1}{2}}\mathcal{A}\Sigma^{-\frac{1}{2}}=\left[\begin{matrix}\Lambda_{1}^{-\frac{1}{2}}A\Lambda_{1}^{-\frac{1}{2}}&\Lambda_{1}^{-\frac{1}{2}}B^{T}\Lambda_{2}^{-\frac{1}{2}}&0\\ -\Lambda_{2}^{-\frac{1}{2}}B\Lambda_{1}^{-\frac{1}{2}}&0&\Lambda_{2}^{-\frac{1}{2}}C^{T}\Lambda_{3}^{-\frac{1}{2}}\\ 0&\Lambda_{3}^{-\frac{1}{2}}C\Lambda_{2}^{-\frac{1}{2}}&0\end{matrix}\right].$ (3.1)
Since the matrices $\Lambda_{1},\Lambda_{2},$ and $\Lambda_{3}$ are SPD, then $\Lambda_{1}^{-\frac{1}{2}}A\Lambda_{1}^{-\frac{1}{2}}$ is SPD, and further $\Lambda_{2}^{-\frac{1}{2}}B\Lambda_{1}^{-\frac{1}{2}}$ and $\Lambda_{3}^{-\frac{1}{2}}C\Lambda_{2}^{-\frac{1}{2}}$ are of full row rank. Consequently, the block structure of $\mathcal{A}$ and the matrix $\Sigma^{-\frac{1}{2}}\mathcal{A}\Sigma^{-\frac{1}{2}}$ are identical. Then by Lemma 3.1, $\mathcal{A}$ is positive stable implies that, the matrix $\Sigma^{-\frac{1}{2}}\mathcal{A}\Sigma^{-\frac{1}{2}}$ and consequently, $\Sigma^{-1}\mathcal{A}$ is positive stable. $\blacksquare$
By [22], the stationary iterative process (2.2) converges to the exact solution of the three-by-three block SPP (1.1) for any initial guess vector if and only if $|\vartheta(\mathcal{T})|<1,$ where $\mathcal{T}$ is the iteration matrix. Now, we establish an if and only if a condition that precisely determines the convergence of the PESS iterative process.
Theorem 3.1.
Let $A\in{\mathbb{R}}^{n\times n}$ be a SPD matrix, and let $B\in{\mathbb{R}}^{m\times n}$ and $C\in{\mathbb{R}}^{p\times m}$ be full row rank matrices, and $s>0$. Then, the PESS iterative process (2.2) converges to the unique solution of the three-by-three block SPP (1.1) if and only if
$(2s-1)|\mu|^{2}+2\Re(\mu)>0,\,\,\forall\,\mu\in\sigma(\Sigma^{-\frac{1}{2}}\mathcal{A}\Sigma^{-\frac{1}{2}}).$ (3.2)
Proof.
From (2.2) we have
$\mathcal{T}={\mathscr{P}}_{PESS}^{-1}\mathcal{Q}_{PESS}=(\Sigma+s\mathcal{A})^{-1}(\Sigma-(1-s)\mathcal{A}),$ (3.3)
where $\Sigma=\mathrm{diag}(\Lambda_{1},\Lambda_{2},\Lambda_{3}).$ Given that $\Lambda_{1},\Lambda_{2}$ and $\Lambda_{3}$ are SPD, then
$\mathcal{T}=\Sigma^{-\frac{1}{2}}(I+s\Sigma^{-\frac{1}{2}}\mathcal{A}\Sigma^{-\frac{1}{2}})^{-1}(I-(1-s)\Sigma^{-\frac{1}{2}}\mathcal{A}\Sigma^{-\frac{1}{2}})\Sigma^{\frac{1}{2}}.$ (3.4)
Thus, (3.4) shows that the iteration matrix $\mathcal{T}$ is similar to $\widetilde{\mathcal{T}},$ where
$\widetilde{\mathcal{T}}=(I+s\Sigma^{-\frac{1}{2}}\mathcal{A}\Sigma^{-\frac{1}{2}})^{-1}(I-(1-s)\Sigma^{-\frac{1}{2}}\mathcal{A}\Sigma^{-\frac{1}{2}}).$
Let $\theta$ and $\mu$ be the eigenvalues of the matrices $\widetilde{\mathcal{T}}$ and $\Sigma^{-\frac{1}{2}}\mathcal{A}\Sigma^{-\frac{1}{2}},$ respectively. Then, it is easy to show that
$\theta=\frac{1-(1-s)\mu}{1+s\mu}.$
Since $\Sigma^{-\frac{1}{2}}\mathcal{A}\Sigma^{-\frac{1}{2}}$ is nonsingular then $\mu\neq 0,$ and we have
$|\theta|=\frac{|1-(1-s)\mu|}{|1+s\mu|}=\sqrt{\frac{(1-(1-s)\Re(\mu))^{2}+(1-s)^{2}\Im(\mu)^{2}}{(1+s\Re(\mu))^{2}+s^{2}\Im(\mu)^{2}}}.$ (3.5)
From (3.5), it is clear that $|\theta|<1$ if and only if
$(1-(1-s)\Re(\mu))^{2}+(1-s)^{2}\Im(\mu)^{2}<(1+s\Re(\mu))^{2}+s^{2}\Im(\mu)^{2}.$
This implies $|\theta|<1$ if and only if $(2s-1)|\mu|^{2}+2\Re(\mu)>0.$ Since, $\vartheta(\mathcal{T})=\vartheta{\widetilde{(}\mathcal{T}}),$ the proof is conclusive. $\blacksquare$
Using the condition in Theorem 3.1, next, we present two sufficient conditions that ensure the convergence of the PESS iterative process. The first one is presented as follows:
Corollary 3.1.
Let $A\in{\mathbb{R}}^{n\times n}$ be a SPD matrix, and let $B\in{\mathbb{R}}^{m\times n}$ and $C\in{\mathbb{R}}^{p\times m}$ be full row rank matrices, and if $s\geq\frac{1}{2},$ then $\vartheta(\mathcal{T})<1,$ i.e., the PESS iterative process always converges to the unique solution of the three-by-three block SPP (1.1).
Proof.
By Lemma 3.2, $\Sigma^{-\frac{1}{2}}\mathcal{A}\Sigma^{-\frac{1}{2}}$ is positive stable, implies that $\Re(\mu)>0$ and $|\mu|>0.$ Thus, the inequality (3.2) holds if $s\geq\frac{1}{2}$. This completes the proof. $\blacksquare$
Remark 3.1.
By applying Corollary 3.1, the unconditional convergence of the existing preconditioners, namely ${\mathscr{P}}_{SS}$, ${\mathscr{P}}_{GSS},$ and ${\mathscr{P}}_{EGSS}$ can be obtained from ${\mathscr{P}}_{PESS}$ by sustituting $s=\frac{1}{2}$.
Next, we present a stronger sufficient condition to Corollary 3.1 for the convergence of the PESS iterative process.
Corollary 3.2.
Let $A\in{\mathbb{R}}^{n\times n}$ be a SPD matrix, and let $B\in{\mathbb{R}}^{m\times n}$ and $C\in{\mathbb{R}}^{p\times m}$ be full row rank matrices, if
$s>\max\left\{\frac{1}{2}\left(1-\frac{\mu_{\min}(\Sigma^{-\frac{1}{2}}(\mathcal{A}+\mathcal{A}^{T})\Sigma^{-\frac{1}{2}})}{\vartheta(\Sigma^{-\frac{1}{2}}\mathcal{A}\Sigma^{-\frac{1}{2}})^{2}}\right),0\right\},$
then the PESS iterative process is convergent for any random initial guess.
Proof.
Note that, as $\Re(\mu)>0,$ the condition in (3.2) holds if and only if $\frac{1}{2}-\frac{\Re(\mu)}{|\mu|^{2}}<s.$ Let ${\bf p}$ be an eigenvector corresponding to $\mu,$ then we have
$\Re(\mu)=\frac{{\bf p}^{*}(\Sigma^{-\frac{1}{2}}(\mathcal{A}+\mathcal{A}^{T})\Sigma^{-\frac{1}{2}}){\bf p}}{2{\bf p}^{*}{\bf p}}\geq\frac{1}{2}\mu_{\min}(\Sigma^{-\frac{1}{2}}(\mathcal{A}+\mathcal{A}^{T})\Sigma^{-\frac{1}{2}}).$
Again, we have $|\mu|\leq\vartheta(\Sigma^{-\frac{1}{2}}\mathcal{A}\Sigma^{-\frac{1}{2}}),$ then we get the following
$\frac{1}{2}-\frac{\mu_{\min}(\Sigma^{-\frac{1}{2}}(\mathcal{A}+\mathcal{A}^{T})\Sigma^{-\frac{1}{2}})}{2\vartheta(\Sigma^{-\frac{1}{2}}\mathcal{A}\Sigma^{-\frac{1}{2}})^{2}}\geq\frac{1}{2}-\frac{\Re(\mu)}{|\mu|^{2}}.$ (3.6)
If we apply $s>0$, then we get the desired result from (3.6). $\blacksquare$
The sufficient condition on Corollary 3.2 is difficult to find for large $\mathcal{A}$ due to the involvement of computation of $\mu_{\min}(\Sigma^{-\frac{1}{2}}(\mathcal{A}+\mathcal{A}^{T})\Sigma^{-\frac{1}{2}})$ and $\vartheta(\Sigma^{-\frac{1}{2}}\mathcal{A}\Sigma^{-\frac{1}{2}})^{2}.$ Thus, we consider $s\geq\frac{1}{2}$ for practical implementation.
4 Spectral distribution of the PESS preconditioned matrix
Solving the three-by-three block SPP (1.1) is equivalent to solving the preconditioned system of linear equations
${\mathscr{P}}_{PESS}^{-1}\mathcal{A}{\bf u}={\mathscr{P}}_{PESS}^{-1}{\bf d},$ (4.1)
where ${\mathscr{P}}_{PESS}$ serves as a preconditioner for the preconditioned GMRES (PGMRES) process. Since preconditioned matrices with clustered spectrum frequently culminate in rapid convergence for GMRES, see, for example [5, 22], the spectral distribution of the preconditioned matrix ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ requires careful attention. As an immediate consequence of Corollary 3.1, the following clustering property for the eigenvalues of ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ can be established.
Theorem 4.1.
Let $A\in{\mathbb{R}}^{n\times n}$ be a SPD matrix, and let $B\in{\mathbb{R}}^{m\times n}$ and $C\in{\mathbb{R}}^{p\times m}$ are full row rank matrices. Suppose that $\lambda$ is an eigenvalue of the PESS preconditioned matrix ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$. Then, for $s\geq\frac{1}{2},$ we have
$|\lambda-1|<1,$
i.e., the spectrum of ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ entirely contained in a disk with centered at $(1,0)$ and radius strictly smaller than $1$
Proof.
Assume that $\lambda$ and $\theta$ are the eigenvalue of the PESS preconditioned matrix ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ and the iteration matrix $\mathcal{T}$, respectively. On the other hand, from the matrix splitting (2.1), we have
${\mathscr{P}}_{PESS}^{-1}\mathcal{A}={\mathscr{P}}_{PESS}^{-1}({\mathscr{P}}_{PESS}-\mathcal{Q}_{PESS})=I-\mathcal{T}.$
Then, we get $\lambda=1-\theta.$ Now, from Corollary 3.1, for $s\geq\frac{1}{2},$ it holds that $|\theta|<1.$ Hence, it follows that, for $s\geq\frac{1}{2},$ $|\lambda-1|<1,$ which completes the proof. $\blacksquare$
Let $(\lambda,{\bf p}=[u^{T},v^{T},w^{T}]^{T})$ be an eigenpair of the preconditioned matrix ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}.$ Then, we have $\mathcal{A}{\bf p}=\lambda{\mathscr{P}}_{PESS}{\bf p},$ which can be written as
$Au+B^{T}v=\lambda(\Lambda_{1}+sA)u+s\lambda B^{T}v,$ (4.2a)
$-Bu-C^{T}w=-s\lambda Bu+\lambda\Lambda_{2}v-s\lambda C^{T}w,$ (4.2b)
$Cv=s\lambda Cv+\lambda\Lambda_{3}w.$ (4.2c)
Remark 4.1.
Notice that from (4.2a)-(4.2c), we get $\lambda\neq\frac{1}{s},$ otherwise , ${\bf p}=[u^{T},v^{T},w^{T}]^{T}={0}\in{\mathbb{R}}^{n+m+p},$ which is impossible as ${\bf p}$ is an eigenvector.
Proposition 4.1.
Let $A\in{\mathbb{R}}^{n\times n}$ be a SPD matrix, and let $B\in{\mathbb{R}}^{m\times n}$ and $C\in{\mathbb{R}}^{p\times m}$ be full row rank matrices, and $s>0.$ Let $(\lambda,{\bf p}=[u^{T},v^{T},w^{T}]^{T})$ be an eigenpair of the preconditioned matrix ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}.$ Then, the following holds
  1. $(1)$
    $u\neq 0.$
  2. $(2)$
    When $\lambda$ is real, $\lambda>0.$
  3. $(3)$
    $\Re(\mu)>0,$ where $\mu=\lambda/(1-s\lambda).$
Proof.
$(1)$ Let $u=0$. Then by (4.2a), we get $(1-s\lambda)B^{T}v=0,$ which shows that $v=0,$ as $B$ has full row rank and $\lambda\neq\frac{1}{s}.$ Furthermore, when combined with (4.2c) leads to $w=0,$ as $\Lambda_{3}$ is SPD. Combining the above facts, it follows that ${\bf p}=[u^{T},v^{T},w^{T}]^{T}=0,$ which is impossible as ${\bf p}$ is an eigenvector. Hence, $u\neq 0.$
$(2)$ Let ${\bf p}=[u^{T},v^{T},w^{T}]^{T}$ be an eigenvector corresponding to the real eigenvalue $\lambda.$ Then, we have ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}{\bf p}=\lambda{\bf p},$ or, $\mathcal{A}{\bf p}=\lambda{\mathscr{P}}_{PESS}{\bf p}.$ Consequently, $\lambda$ can be written as
$\lambda$ $=$ $\frac{{\bf p}^{T}\mathcal{A}{\bf p}+{\bf p}^{T}\mathcal{A}^{T}\bf p}{2({\bf p}^{T}{\mathscr{P}}_{PESS}{\bf p}+{\bf p}^{T}{\mathscr{P}}_{PESS}^{T}\bf p)}$
$=$ $\frac{u^{T}Au}{2(su^{T}Au+u^{T}\Lambda_{1}u+v^{T}\Lambda_{2}v+w^{T}\Lambda_{3}w)}>0.$
The last inequality follows from the assumptions that $A,\Lambda_{1},\Lambda_{2}$, and $\Lambda_{3}$ are SPD matrices.
$(3)$ The system of linear equations (4.2a)-(4.2c) can be reformulated as
$\left\{\begin{array}[]{lcl}Au+B^{T}v=(\lambda/(1-s\lambda))\Lambda_{1}u,&\\ -Bu-C^{T}w=(\lambda/(1-s\lambda))\Lambda_{2}v,&\\ Cv=(\lambda/(1-s\lambda))\Lambda_{3}w.\end{array}\right.$ (4.3)
Now, the systems of linear equations in (4.3) can also be written as $\mathcal{A}{\bf p}=\mu\Sigma{\bf p},$ or, $\Sigma^{-1}\mathcal{A}{\bf p}=\mu{\bf p},$ where $\mu=\lambda/(1-s\lambda).$ By Lemma (3.2), we have $\Sigma^{-1}\mathcal{A}$ is positive stable, which implies $\Re(\mu)>0,$ hence the proof is completed. $\blacksquare$
Remark 4.2.
From Proposition 4.1, it follows that when $\lambda$ is real, then $\mu>0.$ Furthermore, the values of $\lambda$ lies in $(0,1/s)$ for all $s>0$.
In the following theorem, we provide sharper bounds for real eigenvalues of the PESS preconditioned matrix. Before that, we introduce the following notation:
$\xi_{\max}:=\lambda_{\max}(\Lambda_{1}^{-1}A),\leavevmode\nobreak\ \xi_{\min}:=\lambda_{\min}(\Lambda_{1}^{-1}A),\leavevmode\nobreak\ \eta_{\max}:=\lambda_{\max}(\Lambda_{2}^{-1}B\Lambda_{1}^{-1}B^{T}),$ (4.4)
$\leavevmode\nobreak\ \eta_{\min}:=\lambda_{\min}(\Lambda_{2}^{-1}B\Lambda_{1}^{-1}B^{T})\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \theta_{\max}:=\lambda_{\max}(\Lambda_{2}^{-1}C^{T}\Lambda_{3}^{-1}C).$ (4.5)
Theorem 4.2.
Let $A\in{\mathbb{R}}^{n\times n}$ be a SPD matrix, and let $B\in{\mathbb{R}}^{m\times n}$ and $C\in{\mathbb{R}}^{p\times m}$ be full row rank matrices, and $s>0.$ Suppose $\lambda$ is a real eigenvalue of the preconditioned matrix ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}.$ Then
$\lambda\in\left(0,\frac{\xi_{\max}}{1+s\xi_{\max}}\right].$ (4.6)
Proof.
Let ${\bf p}=[u^{T},v^{T},w^{T}]^{T}$ be an eigenvector corresponding to the real eigenvalue $\lambda.$ Then, the system of linear equations in (4.2a)-(4.2c) are satisfied. First, we assume that $v=0.$ Subsequently, (4.2a) and (4.2c) are simplified to
$Au=\lambda(\Lambda_{1}+sA)u,\quad\lambda\Lambda_{3}w=0,$ (4.7)
respectively. The second equation of (4.7) gives $w=0,$ as $\Lambda_{3}$ is SPD. Multiplying $u^{T}$ to the first equation of (4.7) from the left, we get
$\lambda=\frac{u^{T}Au}{u^{T}\Lambda_{1}u+su^{T}Au}=\frac{p}{1+sp},$
where $p={\displaystyle{\frac{u^{T}Au}{u^{T}\Lambda_{1}u}}}>0.$ Now, for any nonzero vector $u\in{\mathbb{R}}^{n},$ we can have
$\lambda_{\min}(\Lambda_{1}^{-\frac{1}{2}}A\Lambda_{1}^{-\frac{1}{2}})\leq\frac{u^{T}Au}{u^{T}\Lambda_{1}u}=\frac{(\Lambda_{1}^{\frac{1}{2}}u)^{T}\Lambda_{1}^{-\frac{1}{2}}A\Lambda_{1}^{-\frac{1}{2}}\Lambda_{1}^{\frac{1}{2}}u}{(\Lambda_{1}^{\frac{1}{2}}u)^{T}\Lambda_{1}^{\frac{1}{2}}u}\leq\lambda_{\max}(\Lambda_{1}^{-\frac{1}{2}}A\Lambda_{1}^{-\frac{1}{2}}).$ (4.8)
Since $\Lambda_{1}^{-\frac{1}{2}}A\Lambda_{1}^{-\frac{1}{2}}$ is similar to $\Lambda_{1}^{-1}A$ and the function $f:{\mathbb{R}}\rightarrow{\mathbb{R}}$ defined by $f(p):=p/(1+sp),$ where $s>0,$ is monotonically increasing in $p>0$, we obtain the following bound for $\lambda:$
$\frac{\xi_{\min}}{1+s\xi_{\min}}\leq\lambda\leq\frac{\xi_{\max}}{1+s\xi_{\max}}.$ (4.9)
[…]
$v=-\frac{1}{\mu}\Lambda_{2}^{-1}Bu.$ (4.10)
Substituting (4.10) in (4.2a) yields
$Au=\lambda(\Lambda_{1}+sA)u+\frac{(1-s\lambda)}{\mu}B^{T}\Lambda_{2}^{-1}Bu.$ (4.11)
Premultiplying both sides of (4.11) by $u^{T},$ we obtain
$\lambda u^{T}Au=\lambda^{2}u^{T}\Lambda_{1}u+s\lambda^{2}u^{T}Au+(1-s\lambda)^{2}u^{T}B^{T}\Lambda_{2}^{-1}Bu,$ (4.12)
which is equivalent to
$\lambda^{2}-\lambda\frac{2sq+p}{s^{2}q+sp+1}+\frac{q}{s^{2}q+sp+1}=0,$ (4.13)
where $q={\displaystyle{\frac{u^{T}B^{T}\Lambda_{2}^{-1}Bu}{u^{T}\Lambda_{1}u}}}.$ By solving (4.13), we obtain the following real solutions for $\lambda$
$\lambda^{\pm}=\frac{2sq+p\pm\sqrt{p^{2}-4q}}{2(s^{2}q+sp+1)}$ (4.14)
where $p^{2}-4q\geq 0.$ Consider the functions $\Phi_{1},\Phi_{2}:{\mathbb{R}}\times{\mathbb{R}}\times{\mathbb{R}}\longrightarrow{\mathbb{R}}$ defined by
$\Phi_{1}(p,q,s)$ $=$ $\frac{2sq+p+\sqrt{p^{2}-4q}}{2(s^{2}q+sp+1)},$
$\Phi_{2}(p,q,s)$ $=$ $\frac{2sq+p-\sqrt{p^{2}-4q}}{2(s^{2}q+sp+1)},$
respectively, where $p,s>0,$ $q\geq 0$ and $p^{2}>4q$. Then $\Phi_{1}$ and $\Phi_{2}$ is monotonically increasing in the argument $p>0,$ and $q\geq 0,$ and decreasing in the argument $q\geq 0$ and $p>0,$ respectively. Now, using (4.8), (4.14) and monotonicity of the functions $\Phi_{1}$ and $\Phi_{2},$ we get the following bounds
$\lambda^{+}=\Phi_{1}(p,q,s)<\Phi_{1}(\lambda_{\max}(\Lambda^{-1}_{1}A),0,s),$ (4.15)
$\Phi_{2}(\lambda_{\max}(\Lambda^{-1}_{1}A),0,s)<\lambda^{-}=\Phi_{2}(p,q,s).$ (4.16)
Combining (4.15) and (4.16), we get the following bounds for $\lambda$
$0<\lambda^{\pm}<\frac{\xi_{\max}}{1+s\xi_{\max}}.$ (4.17)
Next, consider the case $Cv\neq 0.$ Then from (4.2c), we get $w=\frac{1}{\mu}\Lambda_{3}^{-1}Cv.$ Substituting the value of $w$ in (4.2b), we obtain
$(1-s\lambda)Bu+\frac{(1-s\lambda)}{\mu}C^{T}\Lambda_{3}^{-1}Cv=-\lambda\Lambda_{2}v.$ (4.18)
Now, we assume that
$\lambda>\frac{\xi_{\max}}{1+s\xi_{\max}}.$ (4.19)
Since by Remark 4.2, we get $01-s\lambda>$ 0 and the above assertion yields $\mu>\xi_{\max}.$ Hence, the matrix $\mu\Lambda_{1}-A$ is nonsingular and (4.2a) gives
$u=\left(\mu\Lambda_{1}-A\right)^{-1}B^{T}v.$ (4.20)
Substituting (4.20) into (4.18), we obtain
$(1-s\lambda)B\left(\mu\Lambda_{1}-A\right)^{-1}B^{T}v+\frac{(1-s\lambda)}{\mu}C^{T}\Lambda_{3}^{-1}Cv=-\lambda\Lambda_{2}v.$ (4.21)
Note that, $v\neq 0.$ Premultiplying by $v^{T}$ on the both sides of (4.21) leads to the following identity
$(1-s\lambda)v^{T}B\left(\mu\Lambda_{1}-A\right)^{-1}B^{T}v+\lambda v^{T}\Lambda_{2}v=-\frac{(1-s\lambda)}{\mu}v^{T}C^{T}\Lambda_{3}^{-1}Cv.$ (4.22)
On the other hand, $\mu\Lambda_{1}-A>\Lambda_{1}(\mu-\ \lambda_{\max}(\Lambda^{-1}A))I>0,$ and hence a SPD matrix. Consequently, the matrix $B\left(\mu\Lambda_{1}-A\right)^{-1}B^{T}$ is also a SPD matrix and this implies $v^{T}B\left(\mu\Lambda_{1}-A\right)^{-1}B^{T}v>0,$ for all $v\not=0.$ This leads to a contradiction to (4.22) as $v^{T}C^{T}\Lambda_{3}^{-1}Cv\geq 0.$ Hence, the assumption (4.19) is not true and we get
$\lambda\leq\frac{\xi_{\max}}{1+s\xi_{\max}}.$ (4.23)
Again, from ($(1)$) of Proposition 4.1, we have $\lambda>0.$ Combining the above together with (4.9) and (4.17), we obtained the desired bounds in (4.6) for $\lambda.$ Hence, the proof is concluded. ∎
Since the preconditioners ${\mathscr{P}}_{SS}$ and ${\mathscr{P}}_{EGSS}$ are special cases of the PESS preconditioner, next, we obtain refined bounds for real eigenvalues of the SS and EGSS preconditioned matrices from Theorem 4.2.
Corollary 4.1.
Let ${\mathscr{P}}_{SS}$ be defined as in (1.10) with $\alpha>0$ and let $\lambda$ be an real eigenvalue of the SS preconditioned matrix ${\mathscr{P}}_{SS}^{-1}\mathcal{A}.$ Then
$\lambda\in\left(0,\frac{2\kappa_{\max}}{\alpha+\kappa_{\max}}\right],$ (4.24)
where $\kappa_{\max}=\lambda_{\max}(A).$
Proof.
Since ${\mathscr{P}}_{SS}$ is a special case of the PESS preconditioner ${\mathscr{P}}_{PESS}$ for $s=\frac{1}{2}$ as discussed in Table 1, the bounds in (4.24) are obtained by substituting $\Lambda_{1}=\frac{1}{2}\alpha I$ and $s=\frac{1}{2}$ in Theorem 4.2. $\blacksquare$
In the next result, we discuss bounds for the real eigenvalues of the EGSS preconditioned matrix ${\mathscr{P}}_{EGSS}^{-1}\mathcal{A}$.
Corollary 4.2.
Let ${\mathscr{P}}_{EGSS}$ be defined as in (1.11) with $\alpha>0$ and let $\lambda$ be an real eigenvalue of the EGSS preconditioned matrix ${\mathscr{P}}_{EGSS}^{-1}\mathcal{A}.$ Then,
$\lambda\in\left(0,\frac{2\tilde{\kappa}_{\max}}{\alpha+\tilde{\kappa}_{\max}}\right],$ (4.25)
where $\tilde{\kappa}_{\max}=\lambda_{\max}(P^{-1}A).$
Proof.
Since ${\mathscr{P}}_{EGSS}$ is a special case of the PESS preconditioner ${\mathscr{P}}_{PESS}$ for $s=\frac{1}{2}$ as discussed in Table 1, and the bounds are in (4.25) are obtained by substituting $\Lambda_{1}=\frac{1}{2}\alpha P$ and $s=\frac{1}{2}$ in Theorem 4.2. $\blacksquare$
The following result shows the bounds when $\lambda$ is a non-real eigenvalue.
Theorem 4.3.
Let $A\in{\mathbb{R}}^{n\times n}$ be a SPD matrix, and let $B\in{\mathbb{R}}^{m\times n}$ and $C\in{\mathbb{R}}^{p\times m}$ be full row rank matrices. Let $s>0$ and $\lambda$ be a non-real eigenvalue of the preconditioned matrix ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ with $\Im(\lambda)\neq 0.$ Suppose ${\bf p}=[u^{T},v^{T},w^{T}]^{T}$ is the eigenvector corresponding to $\lambda.$ Then the following holds
  1. $(1)$
    If $Cv=0,$ then $\lambda$ satisfies
    $\frac{\xi_{\min}}{2+s\xi_{\min}}\leq|\lambda|\leq\sqrt{\frac{\eta_{\max}}{1+s\xi_{\min}+s^{2}\eta_{\max}}}.$
  2. $(2)$
    If $Cv\neq 0,$ then real and imaginary part of $\lambda/(1-s\lambda)$ satisfies
    $\frac{\xi_{\min}\,\eta_{\min}}{2\left(\xi_{\max}^{2}+\eta_{\max}+\theta_{\max}\right)}\leq\Re\left(\frac{\lambda}{1-s\lambda}\right)\leq\frac{\xi_{\max}}{2}\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \left|\Im\left(\frac{\lambda}{1-s\lambda}\right)\right|\leq\sqrt{\eta_{\max}+\theta_{\max}},$
where $\xi_{\max},\xi_{\min},\eta_{\max},\eta_{\min}$ and $\theta_{\max}$ is defined as in (4.4) and (4.5).
Proof.
$(1)$   Let $\lambda$ be an eigenvalue of ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ with $\Im(\lambda)\neq 0$ and the corresponding eigenvector is ${\bf p}=[u^{T},v^{T},w^{T}]^{T}.$ Then, the system of linear equations in (4.2a)- (4.2c) holds. We assert that, $v\neq 0,$ otherwise from (4.2a), we get $Au=\lambda(\Lambda_{1}+sA)u,$ which further implies $(\Lambda_{1}+sA)^{-\frac{1}{2}}A(\Lambda_{1}+sA)^{-\frac{1}{2}}\bar{u}=\lambda\bar{u},$ where $\bar{u}=(\Lambda_{1}+sA)^{\frac{1}{2}}u.$ Since, $(\Lambda_{1}+sA)^{-\frac{1}{2}}A(\Lambda_{1}+sA)^{-\frac{1}{2}}$ is a SPD matrix and by Lemma 4.1, $u\neq 0,$ and hence $\bar{u}\neq 0,$ which implies $\lambda$ is real, thus a contradiction to $\Im(\lambda)\neq 0.$
Initially, we consider the case $Cv=0.$ The aforementioned discussion suggests that $Au\neq\lambda(\Lambda_{1}+sA)u$ for any $u\neq 0.$ Consequently, $(\mu\Lambda_{1}-A)$ is nonsingular. Again, since $Cv=0,$ from (4.2c), and similar to (4.10)-(4.12), we obtain quadratic equation (4.13) in $\lambda$, which has complex solutions
$\lambda^{\pm i}=\frac{2sq+p\pm i\sqrt{p^{2}-4q}}{2(s^{2}q+sp+1)}$ (4.26)
for $4q>p^{2}.$ From (4.26), we get
$|\lambda^{\pm i}|^{2}=\frac{q}{s^{2}q+sp+1}.$ (4.27)
Now, consider the function $\Psi_{1}:{\mathbb{R}}\times{\mathbb{R}}\times{\mathbb{R}}\longrightarrow{\mathbb{R}}$ defined by
$\Psi_{1}(p,q,s)=\frac{q}{s^{2}q+sp+1},$
where $p,q,s>0.$ Then $\Psi_{1}$ is monotonically increasing in the argument $q>0$ while decreasing in the argument $p>0.$ Due to the fact that $q\leq\lambda_{\max}(\Lambda_{1}^{-1}B^{T}\Lambda_{2}^{-1}B)=\eta_{\max},$ we get the following bounds
$|\lambda^{\pm i}|^{2}=\Psi_{1}(p,q,s)$$\leq$ $\Psi_{1}(\xi_{\min},\eta_{\max},s)=\frac{\eta_{\max}}{1+s\xi_{\min}+s^{2}\eta_{\max}}.$ (4.28)
Again, considering the condition $q>(p^{2}/4),$ (4.27) adheres to the inequality $(p/(2+sp))\leq|\lambda^{\pm i}|.$ Notably, the function $\Psi_{2}:{\mathbb{R}}\longrightarrow{\mathbb{R}}$ defined by $\Psi_{2}(p)=p/(2+sp),$ where $p,s>0,$ is monotonically increasing in the argument $p>0,$ we obtain the following bound
$|\lambda^{\pm i}|=\Psi_{2}(p)\geq\Psi_{2}(\xi_{\min})=\frac{\xi_{\min}}{2+s\xi_{\min}}.$ (4.29)
Hence, combining (4.28) and (4.29), proof for the part $(1)$ follows. $(2)$   Next, consider the case $Cv\neq 0,$ then (4.2c) yields $w=\frac{1}{\mu}\Lambda_{3}^{-1}Cv,$ where $\mu=\lambda/(1-s\lambda).$ Substituting this in (4.2b), we get
$Bu+\frac{1}{\mu}C^{T}\Lambda_{3}^{-1}Cv=-\mu\Lambda_{2}v.$ (4.30)
On the other side, nonsingularity of $(\mu\Lambda_{1}-A)$ and (4.2a) enables us the following identity
$u=(\mu\Lambda_{1}-A)^{-1}B^{T}v.$ (4.31)
Putting (4.31) on (4.30), we get
$B(\mu\Lambda_{1}-A)^{-1}B^{T}v+\frac{1}{\mu}C^{T}\Lambda_{3}^{-1}Cv=-\mu\Lambda_{2}v.$ (4.32)
As $v\neq 0,$ premultiplying by $v^{*}$ on the both sides of (4.32) yields
$v^{*}B(\mu\Lambda_{1}-A)^{-1}B^{T}v+\frac{1}{\mu}v^{*}C^{T}\Lambda_{3}^{-1}Cv=-\mu v^{*}\Lambda_{2}v.$ (4.33)
Consider the following eigenvalue decomposition of $\Lambda_{1}^{-\frac{1}{2}}A\Lambda_{1}^{-\frac{1}{2}}$
$\Lambda_{1}^{-\frac{1}{2}}A\Lambda_{1}^{-\frac{1}{2}}=V\mathcal{D}V^{T},$
where $\mathcal{D}=\mathrm{diag}(\theta_{i})\in{\mathbb{R}}^{n\times n}$ and $\theta_{i}>0,\,i=1,2,\ldots,n$ are in $\sigma(\Lambda_{1}^{-\frac{1}{2}}A\Lambda_{1}^{-\frac{1}{2}})$ and $V\in{\mathbb{R}}^{n\times n}$ is an orthonormal matrix. Then, (4.33) can be written as
$v^{*}B\Lambda_{1}^{-\frac{1}{2}}V(\mu I-\mathcal{D})^{-1}V^{T}\Lambda_{1}^{-\frac{1}{2}}B^{T}v+\frac{1}{\mu}v^{*}C^{T}\Lambda_{3}^{-1}Cv=-\mu v^{*}\Lambda_{2}v.$ (4.34)
Since $(\mu I-\mathcal{D})^{-1}=\Theta_{1}-{\bf{i\mkern 1.0mu}}\Im(\mu)\Theta_{2},$ where ${\bf{i\mkern 1.0mu}}$ denotes the imaginary unit and
$\Theta_{1}={\displaystyle{\mathrm{diag}\left(\frac{\Re(\mu)-\theta_{i}}{(\Re(\mu)-\theta_{i})^{2}+\Im(\mu)^{2}}\right),\Theta_{2}=\mathrm{diag}\left(\frac{1}{(\Re(\mu)-\theta_{i})^{2}+\Im(\mu)^{2}}\right)}},$
from (4.34), and the fact $\Im(\mu)\neq 0,$ we get
$v^{*}B\Lambda_{1}^{-\frac{1}{2}}V\Theta_{1}V^{T}\Lambda_{1}^{-\frac{1}{2}}B^{T}v+\frac{\Re(\mu)}{|\mu|^{2}}v^{*}C^{T}\Lambda_{3}^{-1}Cv=-\Re(\mu)v^{*}\Lambda_{2}v,$ (4.35)
$v^{*}B\Lambda_{1}^{-\frac{1}{2}}V\Theta_{2}V^{T}\Lambda_{1}^{-\frac{1}{2}}B^{T}v+\frac{1}{|\mu|^{2}}v^{*}C^{T}\Lambda_{3}^{-1}Cv=v^{*}\Lambda_{2}v.$ (4.36)
Observed that, $\Theta_{2}\leq\frac{1}{\Im(\mu)^{2}}I,$ then from(4.36), we obtain
$1\leq\frac{1}{\Im(\mu)^{2}}\left(\frac{v^{*}B\Lambda_{1}^{-1}B^{T}v}{v^{*}\Lambda_{2}v}+\frac{v^{*}C^{T}\Lambda_{3}^{-1}Cv}{v^{*}\Lambda_{2}v}\right).$
Therefore, we get
$|\Im(\mu)|\leq\sqrt{\eta_{\max}+\theta_{\max}}.$ (4.37)
Furthermore, notice that $\Theta_{1}\geq\left(\Re(\mu)-\xi_{\max}\right)\Theta_{2},$ then from (4.35), we deduce
$-\Re(\mu)\geq\left(\Re(\mu)-\xi_{\max}\right)\frac{v^{*}B\Lambda_{1}^{-\frac{1}{2}}V\Theta_{2}V^{T}\Lambda_{1}^{-\frac{1}{2}}B^{T}v}{v^{*}\Lambda_{2}v}+\frac{\Re(\mu)}{|\mu|^{2}}\frac{v^{*}C^{T}\Lambda_{3}^{-1}Cv}{v^{*}\Lambda_{2}v}.$ (4.38)
Using (4.36) to (4.38), we get
$2\Re(\mu)\leq\xi_{\max}-\frac{\xi_{\max}}{|\mu|^{2}}\frac{v^{*}C^{T}\Lambda_{3}^{-1}Cv}{v^{*}\Lambda_{2}v}\leq\xi_{\max}.$ (4.39)
Hence, (4.39) yields the following bound
$\Re(\mu)\leq\frac{\xi_{\max}}{2}.$ (4.40)
On the other side, from (4.35) and (4.36), we deduce that
$2\Re(\mu)=\frac{v^{*}B\Lambda_{1}^{-\frac{1}{2}}V(\Re(\mu)\Theta_{2}-\Theta_{1})V^{T}\Lambda_{1}^{-\frac{1}{2}}B^{T}v}{v^{*}\Lambda_{2}v}\geq\xi_{\min}\,\frac{v^{*}B\Lambda_{1}^{-\frac{1}{2}}V\mathcal{X}V^{T}\Lambda_{1}^{-\frac{1}{2}}B^{T}v}{v^{*}\Lambda_{2}v},$ (4.41)
where ${\displaystyle{\mathcal{X}=\mathrm{diag}\left(\frac{1}{(\Re(\mu)-\theta_{i})^{2}+\Im(\mu)^{2}}\right).}}$ Note that
${\displaystyle{\frac{1}{(\Re(\mu)-\theta_{i})^{2}+\Im(\mu)^{2}}\geq\frac{1}{\displaystyle{\max_{i}}(\Re(\mu)-\theta_{i})^{2}+\Im(\mu)^{2}}.}}$
Now, using Proposition 4.1 and (4.40), we get
$-\theta_{i}<\Re(\mu)-\theta_{i}\leq\frac{\xi_{\max}}{2}-\theta_{i},$
which further yields $(\Re(\mu)-\theta_{i})^{2}\leq\max\left\{\left(\frac{\xi_{\max}}{2}-\theta_{i}\right)^{2},\theta_{i}^{2}\right\}.$ Hence, we get
$\displaystyle{\max_{i}}(\Re(\mu)-\theta_{i})^{2}$ $=$ $\displaystyle{\max_{i}}\{(\Re(\mu)-\xi_{\min})^{2},(\Re(\mu)-\xi_{\max})^{2}\}$ (4.42)
$\leq$ $\displaystyle{\max_{i}}\left\{\left(\frac{\xi_{\max}}{2}-\xi_{\min}\right)^{2},\xi^{2}_{\max}\right\}=\xi^{2}_{\max}.$
Therefore, by (4.42) and from the bound in (4.37), we obtain
$\mathcal{X}\geq\frac{I}{\xi^{2}_{\max}+\eta_{\max}+\theta_{\max}}$ (4.43)
Combining (4.41) and (4.43) leads to the following bounds
$2\Re(\mu)$$\geq$ $\frac{\xi_{\min}}{\xi^{2}_{\max}+\eta_{\max}+\theta_{\max}}\frac{v^{*}B\Lambda_{1}^{-1}B^{T}v}{v^{*}\Lambda_{2}v}\geq\frac{\xi_{\min}\eta_{\min}}{\xi^{2}_{\max}+\eta_{\max}+\theta_{\max}}.$ (4.44)
Hence, the proof of part $(2)$ follows by merging the inequalities of (4.37) and (4.44). $\blacksquare$
Remark 4.3.
From the bounds in $(2)$ of Theorem 4.3, we obtain
${\frac{1}{|\tau|^{2}}\left(\frac{\xi_{\min}\eta_{\min}}{2(\xi_{\max}^{2}+\eta_{\max}+\theta_{\max})}+\frac{1}{s}\right)\leq\frac{1}{s}-\Re(\lambda)\leq\frac{1}{|\tau|^{2}}(\frac{\xi_{\max}}{2}+\frac{1}{s})}$
and
$|\Im(\lambda)|\leq\frac{1}{s\tau\bar{\tau}}\sqrt{\eta_{\max}+\theta_{\max}},$
where $\tau=s\mu+1.$ Thus, the real and imaginary parts of the eigenvalues of the preconditioned matrix ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ cluster better as $s$ grows. Therefore, the PESS preconditioner can accelerate the rate of convergence of the Krylov subspace process, like $GMRES$.
Utilizing the established bounds in Theorem 4.3, we derive the subsequent estimations for non-real eigenvalues of SS and EGSS preconditioned matrices.
Corollary 4.3.
Let ${\mathscr{P}}_{SS}$ be defined as in (1.10) with $\alpha>0$ and let $\lambda$ be a non-real eigenvalue of the SS preconditioned matrix ${\mathscr{P}}_{SS}^{-1}\mathcal{A}$ with $\Im(\lambda)\neq 0,$ and ${\bf p}=[u^{T},v^{T},w^{T}]^{T}$ is the corresponding eigenvector. Then,
$(1)$   If $Cv=0,$ $\lambda$ satisfies
$\frac{2\kappa_{\min}}{2\alpha+\kappa_{\min}}\leq|\lambda|\leq\sqrt{\frac{4\tau_{\max}}{\alpha^{2}+\alpha\kappa_{\min}+\tau_{\max}}}.$
$(2)$    If $Cv\neq 0,$ the real and imaginary part of $\lambda/(2-\lambda)$ satisfies
$\frac{\kappa_{\min}\,\tau_{\min}}{2\alpha\left(\kappa^{2}_{\max}+\tau_{\max}+\beta_{\max}\right)}\leq\Re\left(\frac{\lambda}{2-\lambda}\right)\leq\frac{\kappa_{\max}}{2\alpha}\leavevmode\nobreak\ \mbox{and}\leavevmode\nobreak\ \left|\Im\left(\frac{\lambda}{2-\lambda}\right)\right|\leq\frac{\sqrt{\tau_{\max}+\beta_{\max}}}{\alpha},$
where $\kappa_{\min}=\lambda_{\min}(A),$ $\tau_{\max}=\lambda_{\max}(BB^{T}),$ $\tau_{\min}=\lambda_{\min}(BB^{T})$ and $\beta_{\max}=\lambda_{\max}(C^{T}C).$
Proof.
Since ${\mathscr{P}}_{SS}$ is a special case of ${\mathscr{P}}_{PESS}$ for $s=\frac{1}{2}$ as discussed in Table 1, then desired bounds will be obtained by setting $\Lambda_{1}=\Lambda_{2}=\Lambda_{3}=\frac{1}{2}\alpha I$ and $s=\frac{1}{2}$ into Theorem 4.3. $\blacksquare$
Corollary 4.4.
Let ${\mathscr{P}}_{EGSS}$ be defined as in (1.11), and let $P,Q$ and $W$ are SPD matrices with $\alpha,\beta,\gamma>0.$ Let $\lambda$ be a non-real eigenvalue of the EGSS preconditioned matrix ${\mathscr{P}}_{EGSS}^{-1}\mathcal{A}$ with $\Im(\lambda)\neq 0,$ and ${\bf p}=[u^{T},v^{T},w^{T}]^{T}$ is the corresponding eigenvector. Then $(1)$ If $Cv=0,$ $\lambda$ satisfies
$\frac{2\tilde{\kappa}_{\min}}{2\alpha+\tilde{\kappa}_{\min}}\leq|\lambda|\leq\sqrt{\frac{4\tilde{\tau}_{\max}}{\alpha\beta+\beta\tilde{\kappa}_{\min}+\tilde{\tau}_{\max}}}.$
$(2)$   If $Cv\neq 0,$ the real and imaginary part of $\lambda/(2-\lambda)$ satisfies
$\frac{\tilde{\kappa}_{\min}\,\tilde{\tau}_{\min}}{2\left(\beta\tilde{\kappa}^{2}_{\max}+\alpha\tilde{\tau}_{\max}+(\alpha^{2}/\gamma)\tilde{\beta}_{\max}\right)}\leq\Re\left(\frac{\lambda}{2-\lambda}\right)\leq\frac{\tilde{\kappa}_{\max}}{2\alpha}\leavevmode\nobreak\ \mbox{and}$
$\left|\Im\left(\frac{\lambda}{2-\lambda}\right)\right|\leq\sqrt{\frac{1}{\alpha\beta}\tilde{\tau}_{\max}+\frac{1}{\beta\gamma}\tilde{\beta}_{\max}},$
where $\tilde{\kappa}_{\min}=\lambda_{\min}(P^{-1}A),$ $\tilde{\tau}_{\max}=\lambda_{\max}(Q^{-1}BP^{-1}B^{T}),$ $\tilde{\tau}_{\min}=\lambda_{\min}(Q^{-1}BP^{-1}B^{T})$ and $\tilde{\beta}_{\max}=\lambda_{\max}(C^{T}C).$
Proof.
Since ${\mathscr{P}}_{EGSS}$ is a special case of ${\mathscr{P}}_{PESS}$ for $s=\frac{1}{2}$ as discussed in Table 1, then the desired bounds will be obtained by setting $\Lambda_{1}=\frac{1}{2}\alpha P,$ $\Lambda_{2}=\frac{1}{2}\beta Q$, $\Lambda_{3}=\frac{1}{2}\gamma W$ and $s=\frac{1}{2}$ in Theorem 4.3. $\blacksquare$
5 Local PESS (LPESS) preconditioner
To enhance the efficiency of the PESS preconditioner, in this section, we propose a relaxed version of the PESS preconditioner by incorporating the concept of RSS preconditioner [9]. By omitting the term $\Lambda_{1}$ from the $(1,1)$-block of ${\mathscr{P}}_{PESS}$, we present the local PESS (LPESS) preconditioner, denoted as ${\mathscr{P}}_{LPESS}$, defined as follows:
${\mathscr{P}}_{LPESS}:=\left[\begin{matrix}A&sB^{T}&0\\ -sB&\Lambda_{2}&-sC^{T}\\ 0&sC&\Lambda_{3}\end{matrix}\right].$ (5.1)
The implementation of the LPESS preconditioner is similar to the Algorithm 1. However, there is one modification in step $3,$ i.e., we need to solve a linear subsystem $(sA+s^{2}B^{T}\widehat{X}^{-1}B)w_{1}=v$ instead of $\widehat{A}w_{1}=v.$
To illustrate the efficiency of the LPESS preconditioner, we study the spectral distribution of the preconditioned matrix ${\mathscr{P}}_{LPESS}^{-1}\mathcal{A}.$ To achieve this, we consider the following decomposition of the matrices $\mathcal{A}$ and ${\mathscr{P}}_{LPESS}$ as follows:
$\mathcal{A}=\mathfrak{L}\mathfrak{D}\mathfrak{U}\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ {\mathscr{P}}_{LPESS}=\mathfrak{L}\widetilde{\mathfrak{D}}\mathfrak{U},$ (5.2)
where
$\mathfrak{L}=\left[\begin{matrix}I&0&0\\ -BA^{-1}&I&0\\ 0&0&I\end{matrix}\right],\mathfrak{U}=\left[\begin{matrix}I&A^{-1}B^{T}&0\\ 0&I&0\\ 0&0&I\end{matrix}\right],\mathfrak{D}=\left[\begin{matrix}A&0&0\\ 0&Q&-C^{T}\\ 0&C&0\end{matrix}\right],\widetilde{\mathfrak{D}}=\left[\begin{matrix}sA&0&0\\ 0&\Lambda_{2}+sQ&-C^{T}\\ 0&C&\Lambda_{3}\end{matrix}\right],$
and $Q=BA^{-1}B^{T}.$ Using the decomposition in (5.2), we have
${\mathscr{P}}^{-1}_{LPESS}\mathcal{A}=\mathfrak{U}^{-1}\left[\begin{matrix}s^{-1}I&0\\ 0&M^{-1}K\end{matrix}\right]\mathfrak{U},$ (5.3)
where $M=\left[\begin{matrix}\Lambda_{2}+sQ&-sC^{T}\\ sC&\Lambda_{3}\end{matrix}\right]$ and $K=\left[\begin{matrix}Q&-C^{T}\\ C&0\end{matrix}\right].$ Since, ${\mathscr{P}}^{-1}_{LPESS}\mathcal{A}$ is similar to $\left[\begin{matrix}s^{-1}I&0\\ 0&M^{-1}K\end{matrix}\right],$ ${\mathscr{P}}_{LPESS}^{-1}\mathcal{A}$ has eigenvalue $\frac{1}{s}$ with multiplicity $n$ and the remaining eigenvalues satisfies the generalized eigenvalue problem $Kp=\lambda Mp.$
Theorem 5.1.
Let $A\in{\mathbb{R}}^{n\times n},\Lambda_{2}\in{\mathbb{R}}^{m\times m}$ and $\Lambda_{3}\in{\mathbb{R}}^{p\times p}$ be SPD matrices, and let $B\in{\mathbb{R}}^{m\times n}$ and $C\in{\mathbb{R}}^{p\times m}$ be full row rank matrices. Assume that $s>0,$ then LPESS preconditioned matrix ${\mathscr{P}}_{LPESS}^{-1}\mathcal{A}$ has $n$ eigenvalue equal to $\frac{1}{s}.$ The remaining $m+p$ eigenvalues satisfies the following:
  1. $(1)$
    real eigenvalues located in the interval
    $\Big{[}\min\Big{\{}\frac{\vartheta_{\min}}{1+s\vartheta_{\min}},\frac{\theta_{\min}}{\vartheta_{\max}+s\theta_{\min}}\Big{\}},\frac{\vartheta_{\max}}{1+s\vartheta_{\max}}\Big{]}.$
  2. $(2)$
    If $\lambda$ is any non-real eigenvalue (i.e., $\Im(\lambda)\neq 0$), then
    $\frac{\vartheta_{\min}}{2+s\vartheta_{\min}}\leq|\lambda|\leq\sqrt{\frac{\theta_{\max}}{1+s\vartheta_{\min}+s^{2}\theta_{\max}}}\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \frac{1}{s(1+s\sqrt{\theta_{\max}})}\leq|\lambda-\frac{1}{s}|\leq\frac{2}{s(2+s\vartheta_{\min})},$
where $\vartheta_{\max}:=\lambda_{\max}(\Lambda_{2}^{-1}Q),$ $\vartheta_{\min}:=\lambda_{\min}(\Lambda_{2}^{-1}Q)$ and $\theta_{\min}:=\lambda_{\min}(\Lambda_{3}^{-1}C\Lambda_{1}^{-1}C^{T}).$
Proof.
Observe that,
$M^{-1}K=\mathcal{F}^{-1}\widetilde{M}^{-1}\widetilde{K}\mathcal{F},$ (5.4)
where $\mathcal{F}=\left[\begin{matrix}I&0\\ 0&-I\end{matrix}\right],$ $\widetilde{M}=\left[\begin{matrix}\Lambda_{2}+sQ&C^{T}\\ -C&\Lambda_{3}\end{matrix}\right]$ and $\widetilde{K}=\left[\begin{matrix}Q&C^{T}\\ -C&0\end{matrix}\right].$ Hence, $M^{-1}K$ is similar to $\widetilde{M}^{-1}\widetilde{K}.$ Now, $\widetilde{M}^{-1}\widetilde{K}$ is in the form of the preconditioned matrix discussed in [28]. Therefore, by applying the Corollary 4.1 and Theorem 4.3 of [28], we obtain the desired bounds in $(1)$ and $(2).$blacksquare
6 Numerical experiments
In this section, we conduct a few numerical experiments to showcase the superiority and efficiency of the proposed PESS and LPESS preconditioners over the existing preconditioners to enhance the convergence speed of the Krylov subspace iterative process to solve three-by-three block SPPs. Our study involves a comparative analysis among GMRES process and the PGMRES process employing the proposed preconditioners ${\mathscr{P}}_{PESS}$ and ${\mathscr{P}}_{LPESS},$ and the existing preconditioners ${\mathscr{P}}_{BD}$, ${\mathscr{P}}_{IBD}$, ${\mathscr{P}}_{SS},{\mathscr{P}}_{RSS}$ and ${\mathscr{P}}_{EGSS}$ (baseline preconditioners). The numerical results are reported from the aspect of iteration counts (abbreviated as “IT”) and elapsed CPU times in seconds (abbreviated as “CPU”). Each subsystem involving $\widehat{A}$ and $\widehat{X}$ featured in Algorithm 1 is precisely solved by applying the Cholesky factorization of the coefficient matrices. For all iterative process, the initial guess vector is ${\bf u}_{0}=0\in{\mathbb{R}}^{n+m+p}$, and the termination criterion is
${\tt RES}:=\frac{\|\mathcal{A}{\bf u}^{k+1}-{\bf d}\|_{2}}{\|{\bf d}\|_{2}}<10^{-6}.$
The vector ${\bf d}\in{\mathbb{R}}^{n+m+p}$ is chosen so that the exact solution of the system (1.1) is ${\bf u}_{*}=[1,1,\ldots,1]^{T}\in{\mathbb{R}}^{n+m+p}.$ All numerical tests are run in MATLAB (version R2023a) on a Windows 11 operating system with Intel(R) Core(TM) i7-10700 CPU, 2.90GHz, 16 GB memory.
Before proceeding, we discuss the optimal choice for the parameters of PESS preconditioner. We can rewrite the PESS preconditioner as ${\mathscr{P}}_{PESS}=s\widetilde{{\mathscr{P}}}_{PESS},$ where $\widetilde{{\mathscr{P}}}_{PESS}=\left[\begin{matrix}\frac{1}{s}\Lambda_{1}+A&B^{T}&0\\ -B&\frac{1}{s}\Lambda_{2}&-C^{T}\\ 0&C&\frac{1}{s}\Lambda_{3}\end{matrix}\right].$ Since, the pre-factor $s$ has not much effect on the performance of PESS preconditioner, investigating the optimal parameters of ${\mathscr{P}}_{PESS}$ and $\widetilde{{\mathscr{P}}}_{PESS}$ are equivalent. As we know, a preconditioner performs efficiently when it is as close as possible to the coefficient matrix $\mathcal{A},$ observe that $\widetilde{{\mathscr{P}}}_{PESS}-\mathcal{A}={\displaystyle{\frac{1}{s}\Sigma}},$ which tends to a zero matrix as $s$ tends to positive infinity for fixed $\Lambda_{1},\Lambda_{2}$ and $\Lambda_{3}.$ Thus, PESS encounters enhance efficiency for large value of $s.$ However, $s$ can not be too large as the coefficient matrix $\Lambda_{1}+sA+s^{2}B^{T}(\Lambda_{2}+s^{2}C^{T}\Lambda_{3}^{-1}C)^{-1}B$ of the linear subsystem in step $3$ of Algorithm 1 becomes very-ill conditioned. Similar investigations also hold for LPESS preconditioner. Nevertheless, in Figure 7, we show the adaptability of the PESS and LPESS preconditioners by varying the parameter $s.$
Example 1.
We consider the three-by-three block SPP (1.1) taken from [18], where
$A=\left[\begin{matrix}I\otimes G+G\otimes I&0\\ 0&I\otimes G+G\otimes I\end{matrix}\right]\in{\mathbb{R}}^{2l^{2}\times 2l^{2}},\,\,B=\left[\begin{matrix}I\otimes F&F\otimes I\end{matrix}\right]\in{\mathbb{R}}^{l^{2}\times 2l^{2}},$ and $C=E\otimes F\in{\mathbb{R}}^{l^{2}\times l^{2}},$ where $G=\frac{1}{(l+1)^{2}}\,\mathrm{tridiag}(-1,2,-1)\in{\mathbb{R}}^{l\times l},\quad F=\frac{1}{l+1}\,\mathrm{tridiag}(0,1,-1)\in{\mathbb{R}}^{l\times l}$ , and $E=\mathrm{diag}(1,l+1,\ldots,l^{2}-l+1)\in{\mathbb{R}}^{l\times l}.$ Here, $E\otimes F$ denotes the Kronecker product of two matrices $E$ and $F,$ and $\mathrm{tridiag}(a,b,c)\in{\mathbb{R}}^{l\times l}$ denotes the $l\times l$ tridiagonal matrix with diagonal, subdiagonal, and superdiagonal entries are $a,b,$ and $c,$ respectively. For this problem, the size of the matrix $\mathcal{A}$ is $4l^{2}.$
Table 2: Numerical results of GMRES, BD, IBD, SS, RSS, EGSS and PESS PGMRES process for Example 1.
Process $l$ $16$ $32$ $48$ $64$ $80$
size($\mathcal{A}$) $1024$ $4096$ $9216$ $16384$ $25600$
[…]
CPU $0.6607$ $101.7659$ $989.0242$ $2116.1245$ $3755.9251$
RES $078.2852e-$ 07 $079.9189e-$ 07 $079.8389e-$ 07 $031.0813e-$ 03 $031.9862e-$ 03
BD IT $4$ $4$ $4$ $4$ $4$
CPU 0.0611 $2.3135$ $17.0971$ $105.4514$ $337.4322$
RES $131.2728e-$ 13 $131.7577e-$ 13 $136.6653e-$ 13 $126.0211e-$ 12 $122.6610e-$ 12
IBD IT $22$ $22$ $21$ $21$ $21$
CPU 0.0702 $1.1711$ $9.9850$ $44.2197$ $141.1108$
RES $074.0221e-$ 07 $074.6059e-$ 07 $079.1711e-$ 07 $077.8518e-$ 07 $076.8357e-$ 07
Case I
SS IT $4$ $4$ $4$ $4$ $4$
CPU $0.2415$ $1.3375$ $6.7059$ $33.6902$ $111.3118$
RES $087.7528e-$ 08 $085.5120E-$ 08 $084.5134E-$ 08 $083.9157E-$ 08 $083.5073E-$ 08
RSS IT $4$ $4$ $4$ $4$ $4$
CPU $0.2776$ $1.0467$ $6.5207$ $32.4428$ $110.4587$
RES $088.0898e-$ 08 $086.0033e-$ 08 $084.9839e-$ 08 $084.3510e-$ 08 $083.9097e-$ 08
EGSS IT $4$ $4$ $4$ $4$ $4$
CPU $0.3061$ $1.2771$ $6.9552$ $32.8960$ $119.7763$
RES $105.9583e-$ 10 $104.4128e-$ 10 $103.6626e-$ 10 $103.2009e-$ 10 $102.8807e-$ 10
PESS IT 2 2 2 2 2
CPU $0.2285$ $\bf 0.9854$ $\bf 4.8795$ $\bf 23.3687$ $\bf 79.9355$
$s=12$ RES $073.1630e-$ 07 $072.3239e-$ 07 $071.9486e-$ 07 $071.7260e-$ 07 $071.5754e-$ 07
LPESS IT 2 2 2 2 2
CPU $0.2834$ $\bf 0.7989$ $\bf 4.5599$ $\bf 21.1664$ $\bf 72.8376$
$s=12$ RES $073.1180e-$ 07 $072.2463e-$ 07 $071.8481e-$ 07 $071.6071e-$ 07 $071.4411e-$ 07
Case II
SS IT $7$ $7$ $7$ $7$ $7$
CPU 0.2780 $1.6189$ $10.6707$ $54.4493$ $184.6025$
RES $076.7967e-$ 07 $074.9738e-$ 07 $074.1383e-$ 07 $073.6189e-$ 07 $073.2556e-$ 07
RSS IT $7$ $7$ $7$ $7$ $7$
CPU $0.3248$ $1.4644$ $10.4985$ $52.4559$ $177.1895$
RES $075.2397e-$ 07 $073.6832e-$ 07 $072.7532e-$ 07 $072.1869e-$ 07 $071.8240e-$ 07
EGSS IT $5$ $5$ $4$ $4$ $4$
CPU $0.2588$ $1.3792$ $6.7546$ $34.8358$ $118.2298$
RES $086.8491e-$ 08 $071.5239e-$ 07 $079.7099e-$ 07 $076.3466e-$ 07 $074.8159e-$ 07
PESS IT $3$ $3$ $3$ $3$ $3$
CPU $0.2971$ $1.1597$ $6.6682$ $31.2670$ $103.8616$
$s=12$ RES $087.4100e-$ 08 $087.4642e-$ 08 $087.3327e-$ 08 $086.9803e-$ 08 $086.1920e-$ 08
LPESS IT $3$ $3$ $3$ $3$ $3$
CPU $0.2226$ $0.9905$ $6.1339$ $30.0975$ $98.6443$
$s=12$ RES $091.1683e-$ 09 $092.3215e-$ 09 $093.1540e-$ 09 $093.3907e-$ 09 $093.0271e-$ 09
[…]
We consider the parameter choices for the preconditioners ${\mathscr{P}}_{SS},{\mathscr{P}}_{RSS},{\mathscr{P}}_{EGSS},{\mathscr{P}}_{PESS}$ and ${\mathscr{P}}_{LPESS}$ in two cases.
  • In Case I: $\alpha=0.1$ for ${\mathscr{P}}_{SS}$ and ${\mathscr{P}}_{RSS};$ $\alpha=0.1,\beta=1,\gamma=0.001$ and $P=I,Q=I,W=I$ for ${\mathscr{P}}_{EGSS};$ and $\Lambda_{1}=I,\Lambda_{2}=I,\Lambda_{3}=0.001I$ for ${\mathscr{P}}_{PESS}$ and ${\mathscr{P}}_{LPESS}.$
[…]
The parameter selection in Case II is made as in [31]. The optimal value (experimentally) for $s$ in the range $[10,20]$ is determined to be $12$ for minimal CPU times. Following [18], selection of $\widehat{A}$ and $\widehat{S}$ (SPD approximations of $A$ and $S,$ respectively) for IBD preconditioner are done as follows:
$\widehat{A}=LL^{T},\quad\widehat{S}=\mathrm{diag}(B\widehat{A}^{-1}B^{T}),$
where $L$ is the incomplete Cholesky factor of $A$ produced by the Matlab function:
[…]
The test problems generated for the values of $l=16,32,48,64,80,$ and their numerical results are reported in Table 2.
(a) $l=16$
(b) $l=32$
(c) $l=48$
(d) $l=64$
(e) $l=80$
Figure 1: Convergence curves for IT versus RES of PGMRES process employing BD, IBD, SS, RSS, EGSS, PESS and LPESS $(s=12)$ preconditioners in Case II for Example 1.
(a) $\mathcal{A}$
(b) ${\mathscr{P}}_{BD}^{-1}\mathcal{A}$
(c) ${\mathscr{P}}_{IBD}^{-1}\mathcal{A}$
(d) ${\mathscr{P}}_{SS}^{-1}\mathcal{A}$
(e) ${\mathscr{P}}_{RSS}^{-1}\mathcal{A}$
(f) ${\mathscr{P}}_{EGSS}^{-1}\mathcal{A}$
(g) ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ with $s=13$
(h) ${\mathscr{P}}_{LPESS}^{-1}\mathcal{A}$ with $s=13$
Figure 2: Spectral distribution of $\mathcal{A},{\mathscr{P}}_{BD}^{-1}\mathcal{A},{\mathscr{P}}_{SS}^{-1}\mathcal{A},{\mathscr{P}}_{EGSS}^{-1}\mathcal{A},{\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ and ${\mathscr{P}}_{LPESS}^{-1}\mathcal{A}$ for Case II with $l=16$ for Example 1.
Figure 3: Parameter $s$ vs condition number of the preconditioned matrix ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ for Case II with $l=32$ for Example 1.
Figure 1 illustrates convergence curves pertaining to preconditioners BD, IBD, SS, RSS, EGSS, PESS and LPESS ($s=12$) for Case II with $l=16,32,48,64$ and $80$ These curves depict the relationship between the relative residue at each iteration step (RES) and IT counts.
To further illustrate the superiority of the PESS preconditioner, spectral distributions of $\mathcal{A}$ and the preconditioned matrices ${\mathscr{P}}_{BD}^{-1}\mathcal{A},{\mathscr{P}}_{IBD}^{-1}\mathcal{A},{\mathscr{P}}_{SS}^{-1}\mathcal{A},$ ${\mathscr{P}}_{RSS}^{-1}\mathcal{A},{\mathscr{P}}_{EGSS}^{-1}\mathcal{A},{\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ and ${\mathscr{P}}_{LPESS}^{-1}\mathcal{A}$ for the Case II with $l=16$ and $s=13$ are displayed in Figure 2.
To investigate the robustness of the proposed PESS preconditioner, we measure the condition number of the ${\mathscr{P}}_{PESS}^{-1}\mathcal{A},$ which is for any nonsingular matrix $A$ defined by $\kappa(A):=\|A^{-1}\|_{2}\|A\|_{2}.$ In Figure 3, the influence of the parameter $s$ on the condition number of ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ for Case II with $l=32$ is depicted. The parameter $s$ is chosen from the interval $[5,50]$ with step size one. The data presented in Table 2 and Figures 1-3 allow us to make the following noteworthy observations:
  • From Table 2, it is observed that the GMRES has a very slow convergence speed. In both Cases I and II, our proposed PESS and LPESS preconditioners outperform all other compared existing preconditioners from the aspects of IT and CPU times. For example, in Case I with $l=80,$ our proposed PESS preconditioner is $76\%$, $43\%$, $28\%$, $27\%$, and $33\%$ more efficient than the existing BD, IBD, SS, RSS, and EGSS preconditioner, respectively. Moreover, LPESS preconditioners performs approximately $78\%,$ $48\%,$ $35\%,$ $34\%,$ and $39\%$ faster than BD, IBD, SS, RSS, and EGSS preconditioner, respectively. Similar patterns are noticed for $l=16,32,48$ and $64$.
  • In Figure 1, it is evident that the PESS and LPESS preconditioners have a faster convergence speed than the other baseline preconditioners when applied to the PGMRES process for all $l=16,32,48,64$ and $80$
  • According to Figure 2, the spectrum of ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ and ${\mathscr{P}}_{LESS}^{-1}\mathcal{A}$ have better clustering properties than the baseline preconditioned matrices, consequently improves the computational efficiency of our proposed PESS and LPESS PGMRES processes. Moreover, the real eigenvalues of ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ are contained in the interval $(0,0.0714]$ and for ${\mathscr{P}}_{SS}^{-1}\mathcal{A}$ and ${\mathscr{P}}_{EGSS}^{-1}\mathcal{A}$ are in $(0,1.9991]$ and $(0,1],$ respectively, which are consistent with bounds of Theorem 4.2, Corollaries 3.1, and 3.2, respectively. For non-real eigenvalues of ${\mathscr{P}}_{PESS}^{-1}\mathcal{A},$ obtained bounds are consistent with Theorem 4.3.
  • For $l=32,$ the computed condition numbers of $\mathcal{A},{\mathscr{P}}_{BD}^{-1}\mathcal{A},$ and ${\mathscr{P}}_{IBD}^{-1}\mathcal{A}$ are $5.4289e+04,$ $9.5567e+09,$ and $9.5124e+09,$ respectively, which are very large. Whereas Figure 3 illustrate a decreasing trend in the condition number of ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ (for instance, when $s=50$, $\kappa({\mathscr{P}}_{PESS}^{-1}\mathcal{A})=3.4221$) with increasing values of $s.$ This observation highlights that ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ emerges as well-conditioned, consequently the solution obtained by the proposed PESS PGMRES process is more reliable and robust.
The above discussions affirm that the proposed PESS and LPESS preconditioner is efficient, robust, and better well-conditioned with respect to the baseline preconditioners.
Figure 4: Relationship of norm error of solution with increasing noise percentage, employing proposed ${\mathscr{P}}_{PESS}$ for Case II with $s=12$ for $l=16,32,48,64$ and $80$ of Example 1.
Sensitivity analysis of the solution by employing the proposed PESS preconditioner
To study the sensitivity of the solution obtained by employing the proposed PESS PGMRES process to small perturbations on the input data matrices. In the following, we consider the perturbed counterpart $(\mathcal{A}+\Delta\mathcal{A})\tilde{\textbf{u}}=\textbf{d}$ of the three-by-three block SPP (1.1), where the perturbation matrix $\Delta\mathcal{A}$ has same structure as of $\mathcal{A}.$ We construct the perturbation matrix $\Delta\mathcal{A}$ by adding noise to the block matrices $B$ and $C$ of $\mathcal{A}$ as follows:
$\Delta B=10^{-4}*N_{P}*std(B).*randn(m,n)\quad\text{and}\quad\Delta C=10^{-4}*N_{P}*std(C).*randn(p,m),$
where $N_{P}$ is the noise percentage and $std(B)$ and $std(C)$ are the standard deviation of $B$ and $C,$ respectively. Here, $randn(m,n)$ is the is an $m$-by-$n$ normally distributed random matrix, generated by the MATLAB command $randn$. We perform the numerical test for $l=16,32,48,64$ and $80$ for Example 1 using the PESS PGMRES process (Case II, $s=12$). The calculated norm error $\|\tilde{\textbf{u}}-\textbf{u}\|_{2}$ among the solution of the perturbed system and the original system with increasing $N_{P}$ from $5\%$ to $40\%$ with step size $5\%$ are displayed in the figure 4. We noticed that with growing $N_{P},$ the norm error of the solution remains consistently less than $10^{-8},$ which demonstrates the robustness of the proposed PESS preconditioner and the solution u is not sensitive to small perturbation on $\mathcal{A}$.
Example 2.
[…]
$-\Delta{\bf u}+\nabla p=0,\quad\text{in}\quad\Xi,$ (6.1)
$\quad\quad\quad\nabla\cdot{\bf u}=0,\quad\text{in}\quad\Xi.$
[…]
Table 3: Numerical results of GMRES, BD, IBD, SS, RSS, EGSS, PESS and LPESS PGMRES processes for Example 2.
Process h $1/8$ $1/16$ $1/32$ $1/64$
size($\mathcal{A}$) $288$ $1088$ $4224$ $16640$
GMRES IT $158$ $548$ $2048$bf{--}- -
CPU $0.2325$ $0.6960$ $124.7595$ $4545.5072$
RES $078.0113e-$ 07 $078.8023e-$ 07 $079.0887e-$ 07 $064.0481e-$ 06
IBD IT $34$ $35$ $25$ $23$
CPU $\bf 0.0147$ $\bf 0.1166$ $1.9070$ $32.6213$
RES $079.4111e-$ 07 $076.9932e-$ 07 $079.2544e-$ 07 $079.3340e-$ 07
Case I
SS IT $5$ $6$ $9$ $11$
CPU $0.3047$ $0.4422$ $2.4279$ $51.5825$
RES $074.5523e-$ 07 $077.9331e-$ 07 $072.5307e-$ 07 $074.5327e-$ 07
RSS IT $\bf 4$ $\bf 4$ $5$ $5$
CPU $0.2044$ $0.3436$ $1.6047$ $27.6128$
RES $083.4853e-$ 08 $079.2155e-$ 07 $081.7469e-$ 08 $072.7261e-$ 07
EGSS IT $7$ $9$ $11$ $15$
CPU $0.2389$ $0.4051$ $4.4353$ $52.7804$
RES $071.6861e-$ 07 $084.0517e-$ 08 $071.3223e-$ 07 $071.774e-$ 07
PESS IT $4$ $4$ $5$ $5$
CPU $0.2068$ $0.3480$ $1.7633$ $25.2920$
$s=30$ RES $082.3499e-$ 08 $074.2761e-$ 07 $082.1247e-$ 08 $072.0987e-$ 07
LPESS IT $\bf 3$ $\bf{3}$ 4 3
CPU $\bf 0.1990$ $\bf 0.2578$ $\bf 1.4778$ $\bf 17.9730$
$s=30$ RES $071.6606e-$ 07 $078.3803e-$ 07 $104.9289e-$ 10 $075.2658e-$ 07
Case II
SS IT $29$ $38$ $60$ $52$
CPU $0.2999$ $1.1144$ $13.4580$ $213.2381$
RES $077.3560e-$ 07 $079.5790e-$ 07 $079.2515e-$ 07 $079.6425e-$ 07
RSS IT $12$ $13$ $14$ $12$
CPU $0.3797$ $0.6019$ $3.4491$ $50.3476$
RES $073.2456e-$ 07 $075.3086e-$ 07 $076.6600e-$ 07 $074.2834e-$ 07
EGSS IT $9$ $9$ $9$ $7$
CPU $0.2459$ $0.4180$ $2.4320$ $27.9359$
RES $097.9001e-$ 09 $098.6623e-$ 09 $097.4644e-$ 09 $071.2457e-$ 07
PESS IT $5$ $5$ $5$ $5$
CPU $0.2091$ $0.4184$ $1.8118$ $25.8048$
$s=26$ RES $082.3801e-$ 08 $081.1965e-$ 08 $098.4923e-$ 09 $085.9356e-$ 08
LPESS IT $4$ $5$ $\bf 5$ $\bf 5$
CPU $0.2070$ $0.3662$ $\bf 1.5330$ $\bf 23.5997$
$s=26$ RES $072.4250e-$ 07 $128.6182e-$ 12 $115.3930e-$ 11 $108.1253e-$ 10
[…]
(a) $\textbf{h}=1/8$
(b) $\textbf{h}=1/16$
(c) $\textbf{h}=1/32$
(d) $\textbf{h}=1/64$
Figure 5: Convergence curves for IT versus RES of the PGMRES processes by employing IBD, SS, RSS, EGSS, PESS and LPESS preconditioners in Case II for Example 2.
[…]
Table 3 shows that the GMRES process has a very slow convergence speed and also does not converge for $\textbf{h}=1/64$ within $7000$ iterations. The proposed preconditioners require almost five times fewer iterations compared to the IBD preconditioner for convergence. Moreover, in both cases, the PESS and LPESS preconditioners outperform the SS, RSS, and EGSS preconditioners in terms of IT and CPU times. Notably, IT for the PESS preconditioner, in both cases, remains constant, whereas for SS and EGSS, IT increases as the size of $\mathcal{A}$ increases. Furthermore, in Case II with $l=1/64$, our proposed LPESS preconditioner is approximately $28\%$, $90\%$, $53\%$, and $16\%$ more efficient than the existing IBD, SS, RSS, and EGSS preconditioner, respectively. Similar patterns are observed for other values of $l.$
(a) $\mathcal{A}$
(b) ${\mathscr{P}}_{IBD}^{-1}\mathcal{A}.$
(c) ${\mathscr{P}}_{SS}^{-1}\mathcal{A}$
(d) ${\mathscr{P}}_{RSS}^{-1}\mathcal{A}$
(e) ${\mathscr{P}}_{EGSS}^{-1}\mathcal{A}$
(f) ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ with $s=26$
(g) ${\mathscr{P}}_{LPESS}^{-1}\mathcal{A}$ with $s=26$
Figure 6: Spectral distribution of $\mathcal{A},{\mathscr{P}}_{IBD}^{-1}\mathcal{A},{\mathscr{P}}_{SS}^{-1}\mathcal{A},{\mathscr{P}}_{EGSS}^{-1}\mathcal{A},{\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ and ${\mathscr{P}}_{LPESS}^{-1}\mathcal{A}$ for Case II with ${\bf h}=1/8.$
The convergence curves in Figure 5 demonstrate the rapid convergence of the proposed PESS and LPESS PGMRES processes compared to the baseline IBD, SS, RSS, and EGSS ones in terms of RES versus IT counts. Moreover, Figure 6 illustrates the spectral distributions of the original matrix $\mathcal{A},$ and the preconditioned matrices ${\mathscr{P}}_{IBD}^{-1}\mathcal{A},{\mathscr{P}}_{SS}^{-1}\mathcal{A},{\mathscr{P}}_{RSS}^{-1}\mathcal{A},$ ${\mathscr{P}}_{EGSS}^{-1}\mathcal{A},$ ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ and ${\mathscr{P}}_{LPESS}^{-1}\mathcal{A}$ for Case II with ${\bf{h}}=1/8.$ According to Figure 6, eigenvalues of ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ and ${\mathscr{P}}_{LPESS}^{-1}\mathcal{A}$ demonstrate a superior clustering compared to the other preconditioned matrices, which leads to a favorable convergence speed for the proposed PESS and LPESS PGMRES processes.
(a) PESS PGMRES process
(b) LPESS PGMRES process
Figure 7: Characteristic curves for IT of the proposed PESS (left) and LPESS (right) PGMRES processes by varying $s$ in the interval $[1,100]$ with step size $1$ with $\textbf{h}=1/16$ in Case I for Example 2.
Furthermore, the system (1.1) exhibits ill-conditioning nature with $\kappa(\mathcal{A})=5.0701e+05.$ While for Case I with $\mathbf{h}=1/16$, $\kappa({\mathscr{P}}_{PESS}^{-1}\mathcal{A})=1.3353$ and $\kappa({\mathscr{P}}_{LPESS}^{-1}\mathcal{A})=1.2056,$ indicating that the proposed PESS and LPESS preconditioned systems are well-conditioned, ensuring an efficient and robust solution. In addition, to demonstrate the relationship of the parameter $s$ and the convergence speed of the PESS and LPESS preconditioners, we plot graphs of IT counts by varying the parameters $s$ in the interval $[1,100]$ with step size one in Figure 7. We consider $\textbf{h}=1/16$ and other choices for $\Lambda_{1},\Lambda_{2}$ and $\Lambda_{3}$ as in Case I. Figure 7 shows that, with the increasing value of $s,$ decreasing trend in the $IT$ counts for both the PESS and LPESS preconditioner. Moreover, for $s>22,$ using LPESS preconditioner, we can solve the three-by-three $SPP$ mentioned only in $3$ iterations.
7 Conclusion
In this paper, we proposed the PESS iterative process and corresponding PESS preconditioner ${\mathscr{P}}_{PESS}$ and its relaxed variant LPESS preconditioner to solve the three-by-three block SPP (1.1). For the convergence of the proposed PESS iterative process, necessary and sufficient criteria are derived. Moreover, we estimated the spectral bounds of the proposed PESS and LPESS preconditioned matrices. This empowers us to derive spectral bounds for SS and EGSS preconditioned matrices, which have not yet been investigated in the literature to the best of our knowledge. Numerous experimental analyses are performed to demonstrate the effectiveness of our proposed PESS and LPESS preconditioners. The key observations are as follows: $(i)$ the proposed PESS and LPESS preconditioners are found to outperform the existing baseline preconditioners in terms of IT and CPU times. $(ii)$ The proposed preconditioners significantly reduces the condition number of $\mathcal{A}$, consequently showing the its proficiency in solving three-by-three block SPPs. $(iii)$ The PESS and LPESS preconditioned matrices have a more clustered spectral distribution than the baseline preconditioned matrices. $(iv)$ Sensitivity analysis conducted by introducing different percentages of noise on the system (1.1) showcases the robustness of the proposed PESS preconditioner. Further investigation to determine the optimal parameter $s$ can be done in future research endeavors.
Declaration of competing interest
The authors disclose that they have no conflicting interests.
Acknowledgments
Pinki Khatun sincerely acknowledges the Council of Scientific &\&& Industrial Research (CSIR) in New Delhi, India, for financial assistance in the form of a fellowship (File no. 09/1022(0098)/2020-EMR-I).